OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

expand the brackets 2n- 6 + 5=3n-3 2n-1= 3n-3 3n-3= 2n-1 3n-2n=-1+3 n=2

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np